Electronics > Repair

HP 54502A Failing Analogue Trigger

(1/4) > >>

Hi Everyone,
I just got an HP 54502A scope and before anyone jumps at me, yes I have ordered a new dallas NVSRAM chip. Strangely enough this one doesn't seem to have failed(or has it) even though the date code is '89.
I absolutely love how this scope looks and I hope it can be good enough for my home use, I have a nice Rigol DS1054Z at work which I love but it feels like a toy. The HP also has an absolutely massive screen compared to anything I've ever used.
Back to the issue at hand. I have looked everywhere on the internet(including this forum), but I cannot anyone else having an issue with the analogue trigger alone.
Ch1 fails the self test for it, Ch1 passes, but fails the hysteresis test. Also, when cold the DAC fails the self test, but it passes once warmed up.
Anyone know what the failure codes mean(000 000 000 0111 for analogue trig)?
I have managed to connect the external trigger and trigger off of that up to about 10Mhz, after that it was not very stable, but to be fair the signal i was feeding wasn't the best.
This thread seems to be the only thing about this issue and there isn't much to it:

PS: I will attach the service manual here for people looking for it as it was a pain to find (had to use the wayback machine on keysights website).
Still can't find any schematics online.

First check that all power supply voltages are in spec, both for voltage AND ripple!  These series have capacitor leakage problems in the power block.

Then be mindful of the cable you use to run the scope calibration.  It is picky.... found out the trial and error way.

Damn fine scopes as far as the construction goes.  Still hold it's own, except no digital phosphor.

I have played around with the calibration, but I will try again with a better bnc to bnc.
All supply voltages are in spec. I will check ripple too.
It looks like a great scope, I just want it to work. Its the third scope I bought this month, all 3 broken with my luck.
First was a Hitachi V1050-f where everything was pretty much broken(sweep generator, trigger, attenuators on both channels), gave up after replacing all the caps on it(only managed to fix the sweep generator and to an extent the attenuators).
Than I got a Hameg 604 That has some weird vertical shift where the traces are only 1/4 of the way up from center even when level nob is all the way to the top on either channel. Comes and goes which is the worst part.
And then the HP with the trigger.
Good part is that I am learning a lot about analogue circuits and scopes in general. |O

Measured the noise on the PSU, everything around 50mV of high frequency noise, 5.2V line had about 150mV p-p. I couldn't find a table with the tolerances for the rails in the service manual.
Still waiting on the DALLAS chip to arrive, maybe that will fix it.
Any other ideas are very welcomed.

I don't think the NVRAM is going to change anything, when that goes bad it will lose settings when the power is off but as long as the power is on it makes no difference.

Have you tried other cables as someone else suggested? I don't have any experience with HP scopes but Tek scope calibration routines are insanely fussy.


[0] Message Index

[#] Next page

There was an error while thanking
Go to full version